# Random Pareto front surfaces

### v1

@@ 3.1 Order-preserving transformations
The polar parametrisation result in Theorem 3.1 tells us that any Pareto front surface is completely defined by its projected lengths. Conceptually, this means that we can define transformations on the space of Pareto front surfaces by simply applying transformations on the projected lengths. As long as these transformation operations preserve the Pareto partial ordering, then we can be assured that resulting set is also a valid Pareto front surface. In Proposition 3.1, we give two necessary and sufficient conditions for this to happen. The first condition (C1) states that the projected lengths have to be positive—this condition ensures that the resulting Pareto front surface is non-empty. The second condition (C2) is the Pareto optimality condition—this ensures that we cannot find two vectors in the set such that one strongly dominates the other. The necessity of this latter condition follows from the spirit of Lemma 3.1, which presents some equivalent statements for Pareto domination based on the length scalarisation function (2). The proof of both of these result are presented in Section A.2 and Section A.3.
@@
Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@ 3.2 Utility functions
Utility functions are often used in multi-objective optimisation in order to assess the quality of a Pareto front approximation. In this section, we introduce a family of utility functions based on the length scalarisation function (2). We show that this family of utility functions are both interpretable and satisfy many desirable properties such as strict Pareto compliancy (Proposition 3.2).
Length-based R2 utility
@@
$U_{\boldsymbol{\eta},\tau}[Y]:=\mathbb{E}_{\boldsymbol{\lambda}\sim\text{Uniform}(\mathcal{S}_{+}^{M-1})}\biggl{[}\max_{\mathbf{y}\in Y}\tau(s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(\mathbf{y}))\biggr{]}=\mathbb{E}_{\boldsymbol{\lambda}\sim\text{Uniform}(\mathcal{S}_{+}^{M-1})}\bigl{[}\tau(\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[Y]])\bigr{]},$ (6)
for any finite set of objective vectors $Y\subset\mathbb{R}^{M}$, where $\tau:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ is any strictly monotonically increasing transformation that is defined over the space of non-negative scalars. In words, this family of utility functions assesses the quality of a Pareto front approximation based on some notion of the average length away from the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. In this setting, a greater utility is more desirable because it means that the set is much further away, on average, from the reference vector.
@@
Remark 3.5 (Hypervolume scalarisation)
The hypervolume scalarised value can also be interpreted geometrically as being equal to the volume of the hypersphere with diameter $s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(\mathbf{y})\geq 0$, that is $\tau^{\textnormal{HV}}(s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(\mathbf{y}))=\nu[\{\mathbf{z}\in\mathbb{R}^{M}:||\mathbf{z}||_{L^{2}}\leq s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(\mathbf{y})/2\}]$ for any objective vector $\mathbf{y}\in\mathbb{R}^{M}$—see the right of Figure 4 for a visualisation of this interpretation in two dimensions.
@@
By design, the length-based R2 utilities (6) inherit all of the standard properties that any R2 utility satisfies such as the monotone and submodularity property (Tu et al., 2024). In the following, we will additionally show that this special subfamily of utility functions also satisfies the strict Pareto compliancy property over the truncated space (Proposition 3.2). Loosely speaking, the strict Pareto compliancy property (Definition 3.3) states that if a set of vectors is strictly better than another, then it should have a higher utility. This result follows immediately from the fact that the hypervolume indicator is strictly Pareto compliant over the truncated space (Zitzler et al., 2003). We prove this result in Section A.4.
@@ 3.3 Loss functions
Loss functions are an important aspect of decision theory because they quantify the cost associated with observing one element when we expected another. In this section, we define the frontier loss functions, which is a family of loss functions acting on the space of Pareto front surfaces, or more generally the space of polar surfaces $\mathbb{L}_{\boldsymbol{\eta}}$. These loss functions are useful because they give us a way to compare between any two Pareto front surfaces. Later on in Section 4.7, we will show how these loss functions can be used in order to define various probabilistic concepts on the space of random Pareto front surfaces.
@@
for any two polar surfaces $A,B\in\mathbb{L}_{\boldsymbol{\eta}}$, where $S:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ denotes a non-negative scoring function (or loss function). Clearly, different choices of scoring functions will lead to different notions of loss. For example, in Table 1, we present some popular scoring functions which are commonly used in the field of probabilistic forecasting—see Gneiting (2011) for discussion.
@@ 4 Pareto front surface statistics
Real-world problems are typically noisy and subject to many sources of uncertainty. In order to quantify this uncertainty, decision makers often appeal to ideas from probability and statistics. In this section, we will generalise these existing ideas to function over the space of Pareto front surfaces $\mathbb{Y}^{*}_{\boldsymbol{\eta}}$. Precisely, we will study the concept of a random Pareto front surface and showcase how our polar parametrisation result (Theorem 3.1) can be easily used in order to define many statistical quantities of interest. Intuitively, we will treat the Pareto front surface as an infinite-dimensional stochastic process (10) and define the statistics according to its corresponding finite-dimensional distributions. Before we introduce these ideas in more detail, we will first describe the general set up and assumptions that we will be using throughout this section.
@@
Consider a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ denotes the sample space, $\mathcal{F}$ denotes a $\sigma$-algebra and $\mathbb{P}$ denotes a probability measure. In this section, we will be interested in studying the Pareto front surface associated with some vector-valued stochastic process $f:\mathbb{X}\times\Omega\rightarrow\mathbb{R}^{M}$, which we have indexed with the inputs $\mathbf{x}\in\mathbb{X}$. For a fixed and known reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, we denote the corresponding Pareto front surface by
@@
Projected length process
By exploiting the explicit form of the random Pareto front surface in (10), we see that the only dependence on the random quantity $\boldsymbol{\omega}\in\Omega$ arises solely in the projected lengths $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\geq 0$ for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$. Therefore all of the distributional information about the Pareto front surface is completely characterised by its projected length process
@@
To avoid dealing with some unnecessary complications, we will make the following two simplifying assumptions. Firstly, we have Assumption 4.1, which states that the Pareto front surface of interest is non-empty [^5] Note that we could easily lift this non-empty assumption by simply treating every empty Pareto front surface as being equivalent to the degenerate singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$. almost surely. Secondly, we have Assumption 4.2, which states that the stochastic process of interest is bounded almost surely. Together, both of these results ensures that the polar parametrisation of the Pareto front surface, described on the right of (10), holds almost surely. Moreover, Assumption 4.2 also ensures that the projected length $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}]\geq 0$, along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, is bounded almost surely. Consequently, this means that the moments and quantiles of the projected lengths exist and are finite.
@@
The Pareto front surface $Y_{\boldsymbol{\eta},f}^{*}$ is non-empty almost surely.
@@
The stochastic process $f$ is bounded almost surely.
@@
All of the results in this section hold for any input space $\mathbb{X}$. Nevertheless, for computational convenience, most of the examples in this section focus on the practical setting where the input space is finite $|\mathbb{X}|<\infty$. This implies that the projected lengths can be computed exactly because the supremum reduces down to a maximum: $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]=\max_{\mathbf{x}\in\mathbb{X}}s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(f(\mathbf{x},\boldsymbol{\omega}))$ for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$.
@@
Most of the probabilistic concepts which we describe in this section are sensitive to the choice of reference vector (Remark 2.1) and any nonlinear transformations of the objective space (Remark 3.3). For example, in Section 4.1, we define the expected Pareto front surface in the original objective space: $\mathbb{E}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$. But we could have quite easily defined the same expectation in the transformed space, before inverting it back to the original objective space: $\tau^{-1}(\mathbb{E}_{\boldsymbol{\omega}}[Y_{\tau(\boldsymbol{\eta}),\tau\circ f}^{*}(\boldsymbol{\omega})])$, where $\tau:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ is an invertible monotonically increasing function. Clearly, these two quantities are not necessarily equal and therefore some care has to be taken when using one or the other. Note that this issue also arises for the scalar-valued setting as well: $\mathbb{E}[\max_{n}A_{n}]\neq\tau^{-1}(\mathbb{E}[\max_{n}\tau(A_{n})])$ for any collection of scalar-valued random variable $A_{n}$ and any invertible monotonically increasing transformation $\tau:\mathbb{R}\rightarrow\mathbb{R}$.
@@ 4.1 Expectation
Under Assumptions 4.1 and 4.2, the expectation of a Pareto front surface distribution (12) is a Pareto front surface $\mathbb{E}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$.
@@
for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, where $\boldsymbol{\omega}_{n}\in\Omega$ denotes some independent samples of the random parameter. Consequently, we can then define the sample mean Pareto front surface $\hat{\mu}_{\boldsymbol{\eta},N}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ using the projected length estimates given by (13). The fact that this estimated front is also a valid Pareto front surface is a direct consequence of Proposition 4.1. We illustrate an example of the estimated mean Pareto front surface in the left plot of Figure 7.
Remark 4.3 (Bayesian bootstrap)
@@ 4.2 Covariance
for any two positive unit vectors $\boldsymbol{\lambda}_{i},\boldsymbol{\lambda}_{j}\in\mathcal{S}_{+}^{M-1}$, where $\boldsymbol{\omega}_{n}\in\Omega$ denotes some independent samples of the random parameter with $N>1$.
@@ 4.3 Quantiles
We define the quantiles of the Pareto front surface distribution (10) as the collection of marginal quantiles
@@
Under Assumptions 4.1 and 4.2, for any $\alpha\in(0,1)$, the $\alpha$-quantile of a Pareto front surface distribution (15) is a Pareto front surface $\mathcal{Q}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega}),\alpha]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$.
@@
$\hat{q}_{\boldsymbol{\eta},\boldsymbol{\lambda},\alpha,N}:=\mathcal{Q}_{\boldsymbol{\omega}\sim\text{Uniform}(\{\mathbf{w}_{n}\}_{n\leq N})}[\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})],\alpha]$ (16)
for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, where $\boldsymbol{\omega}_{n}\in\Omega$ denotes some independent samples of the random parameter. Consequently, we can then define the empirical quantile Pareto front surface $\hat{q}_{\boldsymbol{\eta},\alpha,N}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ using the projected lengths estimates given by (16). By Proposition 4.2, this empirical Pareto front surface is indeed a valid Pareto front surface. We illustrate an example of this empirical quantile front in the right plot of Figure 7.

### v2

@@ 3.1 Order-preserving transformations
The polar parameterisation result in Theorem 3.1 tells us that any Pareto front surface is completely defined by its projected lengths. Conceptually, this means that we can define transformations on the space of Pareto front surfaces by simply applying transformations on the projected lengths. As long as these transformation operations preserve the Pareto partial ordering, then we can be assured that resulting set is also a valid Pareto front surface. In Proposition 3.1 below, we give two necessary and sufficient conditions for this to happen. The first condition (C1) states that the projected lengths have to be positive—this condition ensures that the resulting Pareto front surface is non-empty. The second condition (C2) is the Pareto optimality condition—this ensures that we cannot find two vectors in the set such that one strongly dominates the other. The necessity of this latter condition follows from the spirit of Lemma 3.1, which presents some equivalent statements for Pareto domination based on the length scalarisation function (2). The proof of both of these result are presented in Section A.3 and Section A.2, respectively.
@@
Consider a polar surface $A\in\mathbb{L}_{\boldsymbol{\eta}}$, then this set is a Pareto front surface, that is $A=\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[A]$, if and only if the following two conditions holds:
@@ 3.2 Utility functions
Utility functions are often used in multi-objective optimisation in order to assess the quality of a Pareto front approximation that is comprised of a finite set of points. In this section, we introduce a family of utility functions based on the length scalarisation function (2). We show that this family of utility functions are both interpretable and satisfy many desirable properties such as strict Pareto compliancy (Proposition 3.2).
Length-based R2 utilities
@@
$\begin{split}U_{\boldsymbol{\eta},\tau}[Y]:=\mathbb{E}_{\boldsymbol{\lambda}\sim\text{Uniform}(\mathcal{S}_{+}^{M-1})}\biggl{[}\max_{\mathbf{y}\in Y}\tau(s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(\mathbf{y}))\biggr{]}\\ =\mathbb{E}_{\boldsymbol{\lambda}\sim\text{Uniform}(\mathcal{S}_{+}^{M-1})}\bigl{[}\tau(\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[\mathcal{Y}_{\boldsymbol{\eta}}^{\text{int}}[Y]])\bigr{]},\end{split}$ (6)
for any finite set of objective vectors $Y\subset\mathbb{R}^{M}$, where $\tau:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$ is any strictly monotonically increasing transformation that is defined over the space of non-negative scalars. In our case, this family of utility functions assesses the quality of a Pareto front approximation based on some notion of the average length away from the reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. In this setting, a greater utility is more desirable because it means that the set is much further away, on average, from the reference vector.
@@
By design, the length-based R2 utilities (6) inherit all of the standard properties that any R2 utility satisfies such as the monotone and submodularity property (Tu et al., 2024, Proposition 3.1). In the following, we will additionally show that this special subfamily of utility functions also satisfies the strict Pareto compliancy property over the truncated space (Proposition 3.2). Loosely speaking, the strict Pareto compliancy property (Definition 3.3) states that if a set of vectors is strictly better than another, then it should have a higher utility. This result follows from the fact that the hypervolume indicator is strictly Pareto compliant over the truncated space (Zitzler et al., 2003). The full proof of this result is presented in Section A.4.
@@ 3.3 Loss functions
Loss functions are an important aspect of decision theory because they quantify the cost associated with performing a particular decision or action. In this section, we define the frontier loss functions, which is a family of loss functions acting on the space of Pareto front surfaces, or more generally the space of polar surfaces $\mathbb{L}_{\boldsymbol{\eta}}$. These loss functions are useful because they give us a way to compare between any two Pareto front surfaces. Later on in Section 4.7, we will show how these loss functions can be used in order to define various probabilistic concepts on the space of random Pareto front surfaces such as the expectation (Section 4.1) and quantiles (Section 4.3).
@@
for any two polar surfaces $A,B\in\mathbb{L}_{\boldsymbol{\eta}}$, where $S:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ denotes a non-negative scoring function (or loss function). Clearly, different choices of scoring functions will lead to different notions of loss. For example, later on in Table 1, we present some popular scoring functions which are commonly used in the field of probabilistic forecasting—see Gneiting (2011) for a discussion.
@@ 4 Pareto front surface statistics
Real-world problems are typically noisy and subject to many sources of uncertainty. In order to quantify this uncertainty, decision makers often appeal to ideas from Probability and Statistics. In this section, we will generalise these existing ideas to function over the space of Pareto front surfaces $\mathbb{Y}^{*}_{\boldsymbol{\eta}}$. Precisely, we will study the concept of a random Pareto front surface and showcase how our polar parameterisation result (Theorem 3.1) can be easily used in order to define many statistical quantities of interest. Intuitively, we will treat the Pareto front surface as an infinite-dimensional random variable (10) and define the statistics according to its corresponding finite-dimensional distributions. Before we introduce these ideas in more detail, we will first describe the general set up and assumptions that we will be using throughout this section.
@@
Consider a standard probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ denotes the sample space, $\mathcal{F}$ denotes a $\sigma$-algebra and $\mathbb{P}$ denotes a probability measure. In this section, we will be interested in studying the Pareto front surface associated with some vector-valued random function $f:\mathbb{X}\times\Omega\rightarrow\mathbb{R}^{M}$, which we have indexed with the inputs $\mathbf{x}\in\mathbb{X}$. For a fixed and known reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$, we denote the corresponding Pareto front surface by
@@
Projected lengths
By exploiting the explicit form of the random Pareto front surface in (10), we see that the only dependence on $\boldsymbol{\omega}\in\Omega$ arises solely in the projected lengths $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\geq 0$ for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$. Therefore all of the distributional information about the Pareto front surface is completely characterised by its projected lengths
@@
To avoid dealing with some unnecessary complications, we will make the following two simplifying assumptions (footnote 5 and Assumption 4.2). Together, both of these assumptions ensures that the polar parameterisation of the Pareto front surface, described on the right of (10), holds almost surely. Moreover, Assumption 4.2 also ensures that the projected length $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}]\geq 0$, along any positive direction $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, is bounded almost surely. Consequently, this means that the moments and quantiles of the projected lengths exist and are finite.
@@
The Pareto front surface $Y_{\boldsymbol{\eta},f}^{*}$ is non-empty [^5] Note that we could easily lift this non-empty assumption by simply treating every empty Pareto front surface as being equivalent to the degenerate singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$. almost surely.
@@
The random function $f$ is bounded almost surely.
@@
All of the results in this section hold for any input space $\mathbb{X}$. Nevertheless, for computational convenience, most of the examples in this section focus on the practical setting where the input space is finite and $|\mathbb{X}|<\infty$. This implies that the projected lengths can be computed exactly because the supremum reduces down to a maximum: $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]=\max_{\mathbf{x}\in\mathbb{X}}s_{\boldsymbol{\eta},\boldsymbol{\lambda}}(f(\mathbf{x},\boldsymbol{\omega}))$ for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$.
@@
Most of the probabilistic concepts which we describe in this section depend on the choice of reference vector (Remark 2.1) and any nonlinear transformations of the objective space (Remark 3.3). For example, in Section 4.1, we define the expected Pareto front surface in the original objective space: $\mathbb{E}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$. But we could have quite easily defined the same expectation in the transformed space, before inverting it back to the original objective space: $\tau^{-1}(\mathbb{E}_{\boldsymbol{\omega}}[Y_{\tau(\boldsymbol{\eta}),\tau\circ f}^{*}(\boldsymbol{\omega})])$, where $\tau:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ is an invertible monotonically increasing function. Clearly these two quantities are not necessarily equal and therefore some care has to be taken when using one or the other. Note that this issue also arises for the scalar-valued setting as well.
@@ 4.1 Expectation
Under footnotes 5 and 4.2, the expectation of a Pareto front surface distribution (12) is a Pareto front surface $\mathbb{E}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$.
@@
for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, where $\{\boldsymbol{\omega}_{n}\}_{n=1}^{N}\subset\Omega$ denotes i.i.d. samples of the random parameter. Consequently, we can then define the sample mean Pareto front surface $\hat{\mu}_{\boldsymbol{\eta},N}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ using the projected length estimates given by (13). The fact that this estimated front is also a valid Pareto front surface is a direct consequence of Proposition 4.1. We illustrate an example of the estimated mean Pareto front surface in the left plot of Figure 7.
Example 4.1 (Bayesian bootstrap)
@@ 4.2 Covariance
for any two positive unit vectors $\boldsymbol{\lambda}_{i},\boldsymbol{\lambda}_{j}\in\mathcal{S}_{+}^{M-1}$, where $\{\boldsymbol{\omega}_{n}\}_{n=1}^{N}\subset\Omega$ denotes i.i.d. samples of the random parameter with $N>1$.
@@ 4.3 Quantiles
We also define the quantiles of the Pareto front surface distribution (10) as the collection of marginal quantiles
@@
Under footnotes 5 and 4.2, for any $\alpha\in(0,1)$, the $\alpha$-quantile of a Pareto front surface distribution (15) is a Pareto front surface $\mathcal{Q}_{\boldsymbol{\omega}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega}),\alpha]\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$.
@@
$\hat{q}_{\boldsymbol{\eta},\boldsymbol{\lambda},\alpha,N}:=\mathcal{Q}_{\boldsymbol{\omega}\sim\text{Uniform}(\{\boldsymbol{\omega}_{n}\}_{n=1}^{N})}[\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\omega})],\alpha]$ (16)
for $\boldsymbol{\lambda}\in\mathcal{S}_{+}^{M-1}$, where $\{\boldsymbol{\omega}_{n}\}_{n=1}^{N}\subset\Omega$ denotes i.i.d. samples of the random parameter. Consequently, we can then define the empirical quantile Pareto front surface $\hat{q}_{\boldsymbol{\eta},\alpha,N}\in\mathbb{Y}^{*}_{\boldsymbol{\eta}}$ using the projected lengths estimates given by (16). By Proposition 4.2, this empirical Pareto front surface is indeed a valid Pareto front surface. We illustrate an example of this empirical quantile front in the right plot of Figure 7.